OpenStudy (cheesenips111111):

The incomplete work of a student to solve an equation is shown below: Step 1: 4x + 12 = 4 Step 2: ? Step 3: x = −8 ÷ 4 Step 4: x = −2 What is the missing Step 2?

OpenStudy (cheesenips111111):

i got d

OpenStudy (cheesenips111111):

4x = 8 4x = 16 4x = −16 4x = −8

TheSmartOne (thesmartone):

That is correct. Great work! :)

OpenStudy (cheesenips111111):

can you check more?

OpenStudy (cheesenips111111):

What is the value of z for the equation fraction 1 over 3z = −fraction 5 over 6 + fraction 1 over 6z? 5 2 −2 −5

OpenStudy (cheesenips111111):

i got d

TheSmartOne (thesmartone):

I've seen that question so many times now. xD Your answer is correct :)

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

A system of equations is shown below: Equation A: 2c = d − 8 Equation B: c = 3d + 8 Which of the following steps should be performed to eliminate variable d first? Multiply equation B by 2. Multiply equation A by 2. Multiply equation B by 3. Multiply equation A by −3.

OpenStudy (cheesenips111111):

d again

TheSmartOne (thesmartone):

Correct :)

OpenStudy (cheesenips111111):

What is the value of z in the equation 2z + 6 = −4? 5 1 −1 −5

OpenStudy (cheesenips111111):

i got b

OpenStudy (cheesenips111111):

@TheSmartOne

TheSmartOne (thesmartone):

Nope 2z + 6 = −4 Subtract 6 on both sides Then divide by 2 on both sides

OpenStudy (cheesenips111111):

-1?

OpenStudy (cheesenips111111):

@TheSmartOne

TheSmartOne (thesmartone):

You're adding 6 on both sides, it seems 2z + 6 = −4 Subtract 6 on both sides 2z + 6 - 6 = -4 - 6 2z = -4 - 6 What is -4 - 6 = ? Hint: -4 - 6 = -1(4 + 6) = ??

OpenStudy (cheesenips111111):

idk

OpenStudy (cheesenips111111):

oh wait

OpenStudy (cheesenips111111):

is it 5?

OpenStudy (cheesenips111111):

??

OpenStudy (cheesenips111111):

you there????????????????????????????????????????/

OpenStudy (cheesenips111111):

ill open a new one